• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name olivier

    Log entry time 22:07:42 on October 5,2002

    Entry number 87395

    This entry is a followup to: 87390


    keyword=optics target unavailable

    We have the same problem than last night that the ion chamber in the beam dump tripped at very low beam current when the optics target is in place. Following Dave Meekins' instructions, we moved this target to different positions, but it didn't correct the problem. The next hypothesis is that something happened on this target when the deuterium cell ruptured.

    Anyway the conclusion is that the optics target is unusable at the time, and we are not going to fix it tonight or tomorrow. We will just drop the optics measurements (at least for now).